a line with equation y=5+7x has what slope ?
Found 3 solutions by richard1234, edjones, ewatrrr:
Answer by richard1234(7193)

Slope is 7. In general, if a linear equation is in the form y = mx + b, then m is the slope.

Answer by edjones(8007)

7, the number multiplied times x in the slope-intercept form (which this equation is in.)

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
y = 7x + 5 Slope m = 7
Points (0,5), (-5/7,0) and (-1,-2) on the line